OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 16 + 10x = 14x − 4 is given below. Choose the step that is missing from the solution. 16 + 10x = 14x − 4 16 = 4x − 4 ? 5 = x

OpenStudy (anonymous):

Remember you can group numbers and terms together and that you can move things to the other side of the equation sign by changing its sign.

OpenStudy (ash2326):

16=4x-4 add 4 to both the sides @DB9898

OpenStudy (anonymous):

goku wins agan

OpenStudy (andriod09):

not true ash you subtract for from both sides

OpenStudy (ash2326):

\[16=4x-4\] \[16+4=4x-4+4\]

OpenStudy (ash2326):

@andriod09 there you see, I have added 4 to both the sides

OpenStudy (ash2326):

@DB9898 could you take it from here?

OpenStudy (andriod09):

16+10x=14x-4 -04 = -4 14+10x=14x -14 = -14 10x=14x-14